OpenStudy (anonymous):

use the formula d=rt find the value of the missing variable. d=100 miles,r=50 mph

OpenStudy (anonymous):

100=50t so solve for t

OpenStudy (zzr0ck3r):

d=rt implies t=d/r implies r = d/t

OpenStudy (anonymous):

t=___hr